OpenStudy (anonymous):

Please help! given f(x)=x^2+5 find f(x+h)-f(x)/h

OpenStudy (solomonzelman):

\(f(x+h)=?\)

OpenStudy (solomonzelman):

(Plug in "x+h" instead of "x" into the function, and simplify)

OpenStudy (tkhunny):

Let's be more careful with notation. You have written: \(f(x+h) - \dfrac{f(x)}{h}\) You probably meant [f(x_h) - f(x)]/h = \(\dfrac{f(x+h) - f(x)}{h}\)

OpenStudy (anonymous):

tkhunny yes i meant that

OpenStudy (tkhunny):

You have to write it if you mean it. :-)